OpenStudy (anonymous):

I have this problem and I have to solve for x.. 4x+3y=8 Is the answer x+-3/4y+2 ??? I need to know for my test!!!

OpenStudy (anonymous):

I mean x=-3/4y+2

OpenStudy (anonymous):

Isolate Y and then solve for it, after you have Y's value plug it in and solve for X, system of equations suck

OpenStudy (diyadiya):

x=(8-3y)/4

OpenStudy (anonymous):

So is my answer correct Jshooz?

OpenStudy (anonymous):

Diyadiya is right

OpenStudy (anonymous):

That's none of the choices though?

OpenStudy (anonymous):

Can u type out the entire question then>

OpenStudy (anonymous):

4x+3y=8

OpenStudy (anonymous):

Answers are: A.) x=3/4y+2 B.)x=-3/4y+1/2 C.)4/3y+1/2 D.)-3/4y+2

OpenStudy (anonymous):

C and D are also X=

OpenStudy (anonymous):

D is the right answer

OpenStudy (anonymous):

THANKYOU

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

what about -5x+3y=-9

OpenStudy (anonymous):

I need to get y